# Explainable Machine Learning-Based Security and Privacy Protection Framework for Internet of Medical Things Systems

### v2

@@ 3.4 XAI Process
ML has demonstrated its effectiveness in anomaly detection, particularly in identifying zero-day attacks and new vulnerabilities. This makes ML more advantageous compared to signature-based methods. However, non-transparent ML models, such as ANN, operate as black boxes, making it challenging to explain the reasoning behind the classification of instances, such as identifying an attack. This results in time consumption and ambiguity in analyzing predictions generated by these models. Explainable AI provides a solution by assisting model designers in determining the impact of each attribute on the classification process. This proves beneficial for model designers, aiding in debugging, enhancing the ML model, gaining insights into its decision-making process, and fostering increased trust in the model.
The implementation of ANN involves deploying a non-transparent model, necessitating post-hoc explanation. Consequently, the feature relevance method was employed, as it aligns well with the framework. Among the various methodologies for determining feature relevance, Shapley Additive Explanations (SHAP) stands out as a widely utilized and agnostic method, introduced by Lundberg and Lee [54]. Rooted in cooperative game theory, SHAP relies on Shapley values, offering both local and global scopes of explanation. Within the proposed framework, the global scope explanation has been applied.
@@
$\phi_{i}(v):\text{Shapley value for player (feature) }i\text{ in the cooperative game }v.$
@@
$v(S\cup\{i\}):\text{The value of the coalition }S\text{ extendedby adding player }i.$
@@ 4.2 Dataset Description
Experiments are carried out on four distinct datasets containing attacks that can potentially impact the availability, confidentiality, and integrity of IoMT systems. These datasets include NSL-KDD, UNSW-NB15, and ToN-IoT, all comprising network data, and WUSTL-EHMS, which encompasses both network and medical data. Each dataset is described in detail as follows:
  • NSL-KDD : The KDD99 and NSL-KDD datasets [15] were created by the IST division at the Lincoln Laboratories of the Massachusetts Institute of Technology. To generate the DARPA 98 dataset from raw packets, they developed a simulation testbed within the U.S. Air Force LAN system that included both normal and attack scenario traffic. Subsequently, this dataset was renamed KDD99 and included data characteristics derived from packets. The NSL-KDD dataset, an improved version of KDD99, was later developed to address limitations of the original dataset, such as removing redundant data and achieving a better balance between samples in the training and testing sets. The NSL-KDD dataset encompasses various attack categories, including Probing, Remote to Local (R2L), DoS, and User to Root (U2R), alongside a ”Normal” class representing legitimate network traffic. Comprising 41 features, the dataset includes network connection attributes such as protocol type, service, source and destination IP addresses, and source and destination ports, among others.
  • UNSW-NB15 : The UNSW-NB15 dataset was officially released in 2015 by the Cyber Range Lab [28], which operates under the auspices of the prestigious Australian Center for Cyber Security. Due to its remarkable utility, the dataset has become a common choice for researchers within the cyber security domain, particularly among the research community affiliated with the Australian Centre for Cyber Security (ACCS). In the case of the UNSW-NB15 dataset, the authors opted to utilize unprocessed network packets, which were generated using the highly regarded IXIA Perfectstorm program. As part of the dataset evaluation, a comprehensive range of nine attack scenarios were meticulously implemented, encompassing diverse types such as DoS, fuzzes, analysis, backdoor, generic, reconnaissance, shellcode, exploits, and worms. To provide a holistic representation of network traffic, the dataset also includes a dedicated ”Normal” class, specifically designed to capture legitimate network activity. Notably, a total of 49 network traffic features were meticulously extracted from the dataset, employing the robust Argus and Bro-IDS programs as essential analytical tools.
@@
  • WUSTL-EHMS : Using a real-time Enhanced Healthcare Monitoring System (EHMS) testbed, the WUSTL-EHMS dataset was generated [17]. Due to the limited availability of a dataset that integrates these biometrics, this testbed accrues both the network flow metrics and patients’ biometrics. This dataset comprises MITM attacks, spoofing and data injection. The spoofing attack merely sniffs the protocols passing through the gateway and the server, thus infringing upon the confidentiality of the patient’s data. The data injection attack is employed to dynamically modify the packets, thereby infringing upon the integrity of the data.
@@ 4.3 Data Preprocessing
During the data preprocessing phase, a diverse array of methods is employed to optimize the dataset for analysis. To begin, a mapping function is utilized to assign binary values, effectively distinguishing instances representing attacks from those that do not, thereby enabling a clear classification of the data. The transformation of Boolean values into binary further streamlines the dataset’s representation. For categorical features, the ordinal and OneHotEncoder techniques are applied, ensuring their proper handling in the subsequent analysis. To maintain an unbiased approach, certain features primarily used for dataset labeling are meticulously removed. Additionally, features that hold a single value, bearing no meaningful information for the model’s learning process, are excluded from the dataset. Subsequently, numerical features undergo standardization using the StandardScaler method, optimizing their scale and improving their compatibility with the model.
Ultimately, the dataset is partitioned into an 80% training set and a 20% testing set. In the context of FL, the 80% of the training data are independently and identically distributed among clients, while the remaining 20% of the test set is utilized for evaluating the global model. This division facilitates the model’s learning process on the training set, while the testing set provides an independent evaluation of the model’s performance. These preprocessing steps are essential in preparing the data for effective model training and evaluation.
@@ 4.4 ML Algorithm
The construction of various ANN models tailored to specific datasets involves distinctive configurations. For the UNSW-NB15 dataset, ANNs are constructed with seven hidden layers, each comprising a different number of units: 150, 120, 90, 60, 30, 20, and 10 units, respectively. Conversely, the ToN-IoT dataset utilizes an ANN model with five hidden layers, incorporating 60, 40, 30, 20, and 10 units, respectively. In the case of the NSL-KDD dataset, a five-layered ANN model is established, featuring 80, 40, 30, 20, and 10 units. An ANN model is constructed for the WUSTL-EHMS dataset, comprising three hidden layers that are specified to contain 10, 20, and 40 units, respectively.
@@ 4.5 Evaluation Metrics
  • Loss : is a mathematical function used to quantify the disparity between the anticipated output of a ML model and the factual target value associated with a certain input [61].
@@ 4.6.1 FL Parameters Selections
To evaluate the impact of varying the number of clients on performance and communication cycles, the local epochs are fixed at 1, the fraction fit at 1, and the number of clients is adjusted to 2, 4, 8, and 12. Increasing the number of clients results in a significant reduction in communication cycles required to achieve target performance levels. This trend is evident in the Ton-IoT and NSL-KDD datasets, where the lowest number of communication cycles is achieved with 12 clients, and no significant variation is observed with 8 clients compared to other configurations, as shown in Tables 3(a) and 3(c). This suggests that communication cycles stabilize with 8 or more clients. For the UNSW_NB15 dataset, the lowest number of communication cycles is achieved with 12 clients, showing a significant reduction compared to 8 clients, as indicated in Table 3(b). This aligns with the decreasing trend in communication cycles as the number of clients increases, consistent with observations from the Ton-IoT and NSL-KDD datasets. In contrast, the WUSTL-EHMS dataset exhibits an inverse trend, where increasing the number of clients leads to a higher number of communication rounds to achieve target performance, as depicted in Table 3(d). This behavior is likely due to the limited sample size in the dataset, resulting in insufficient data distribution across clients and requiring additional communication rounds for model generalization.
@@
Increasing the fraction fit effectively reduces communication cycles while maintaining target performance levels. This trend is observed in the Ton-IoT and NSL-KDD datasets, as illustrated in Tables 4(a) and 4(c). However, for the UNSW-NB15 and WUSTL-EHMS datasets, the lowest number of communication rounds is achieved with a fraction fit of 0.5, with a slight increase observed at a fraction fit of 1, as shown in Tables 4(b) and 4(d). This confirms the general trend of reduced communication rounds with higher fraction fit values.
Finally, the analysis investigates the impact of the number of local epochs on performance and communication cycles. The fraction fit is set to 1, as this configuration has been shown to minimize communication rounds. The number of local epochs is varied to 1, 2, 5, and 8.
@@
A notable finding is the consistent reduction in communication cycles with an increase in local epochs while maintaining strong performance. This trend is observed across most datasets, as shown in Tables 5(a), 5(b), and 5(d). However, the NSL-KDD dataset achieves the optimal number of communication rounds with only 2 epochs, as illustrated in Table 5(c), though the results remain close to those obtained with 8 epochs. This further supports the overall trend of reduced communication rounds with increased local epochs.
@@ 4.6.3 Centralized vs FL :
The results are consolidated into stacked bar plots, as depicted in Figure 8, where FL results are represented in orange, and centralized results are represented in blue. The outcomes on the NSL-KDD, TON-IOT, UNSW_NB15, and WUSTL-EHMS datasets, presented in Tables 8(a), 8(b), 8(c), and 8(d) respectively, show bars being halved, indicating comparable results between FL and centralized approaches. Consequently, training models on separate data partitions generalizes as effectively as training a global model on the entire dataset. This dual advantage of achieving comparable detection outcomes while preserving data privacy highlights the efficacy and privacy-centric nature of the FL-based IDS, positioning it as a robust and privacy-aware solution for intrusion detection.
@@ 6 DISCUSSION
The proposed solution places a paramount emphasis on safeguarding patient privacy by employing FL, which involves sharing model weights instead of actual patient data. By optimizing FL parameters, the communication rounds was significantly reduced, thereby minimizing bandwidth consumption, preventing network congestion, and ensuring the scalability of the system. The distributed nature of FL proves instrumental in positioning the IDS in close proximity to potential attack sources. This proximity enhances the system’s agility, allowing for rapid and effective detection and response to security threats. This decentralized approach strengthens the overall security posture while contributing to the rapid identification and mitigation of potential risks to patient data within the IoMT ecosystem.

### v3

@@ 3.4 XAI Process
ML has demonstrated its effectiveness in anomaly detection, particularly in identifying zero-day attacks and new vulnerabilities. This makes ML more advantageous compared to signature-based methods. However, non-transparent ML models, such as ANN, operate as black boxes, making it challenging to explain the reasoning behind the classification of instances, such as identifying an attack. This results in time consumption and ambiguity in analyzing predictions generated by these models. XAI provides a solution by assisting model designers in determining the impact of each attribute on the classification process. This proves beneficial for model designers, aiding in debugging, enhancing the ML model, gaining insights into its decision-making process, and fostering increased trust in the model.
The implementation of ANN involves deploying a non-transparent model, necessitating post-hoc explanation. Consequently, the feature relevance method is employed, as it aligns well with the framework. Among the various methodologies for determining feature relevance, Shapley Additive Explanations (SHAP) stands out as a widely utilized and agnostic method, introduced by Lundberg and Lee [52]. Rooted in cooperative game theory, SHAP relies on Shapley values, offering both local and global scopes of explanation. Within the proposed framework, the global scope explanation is applied.
@@
[…]
@@ 4.2 Dataset Description
Experiments are carried out on four distinct datasets containing attacks that can potentially impact the availability, confidentiality, and integrity of IoMT systems. These datasets include NSL-KDD, UNSW-NB15, and ToN-IoT, all comprising network data, and WUSTL-EHMS, which encompasses both network and medical data.
Each dataset exhibits unique characteristics in terms of attack diversity, feature representation, and data composition, as detailed in the subsequent descriptions. Figure 4 illustrates the comparative distribution of normal versus anomalous instances across all datasets, providing a comprehensive overview of their respective data distributions.
Figure 4: Class Distribution of Normal vs. Anomalous Instances Across Datasets
  • NSL-KDD : The KDD99 and NSL-KDD datasets [14] are created by the IST division at the Lincoln Laboratories of the Massachusetts Institute of Technology. To generate the DARPA 98 dataset from raw packets, they develop a simulation testbed within the U.S. Air Force LAN system that includes both normal and attack scenario traffic. Subsequently, this dataset is renamed KDD99 and includes data characteristics derived from packets. The NSL-KDD dataset, an improved version of KDD99, is later developed to address limitations of the original dataset, such as removing redundant data and achieving a better balance between samples in the training and testing sets. The NSL-KDD dataset encompasses various attack categories, including Probing, Remote to Local (R2L), DoS, and User to Root (U2R), alongside a ”Normal” class representing legitimate network traffic. Comprising 41 features, the dataset includes network connection attributes such as protocol type, service, source and destination IP addresses, and source and destination ports, among others.
  • UNSW-NB15 : The UNSW-NB15 dataset is officially released in 2015 by the Cyber Range Lab [26], which operates under the auspices of the prestigious Australian Center for Cyber Security. Due to its remarkable utility, the dataset becomes a common choice for researchers within the cyber security domain, particularly among the research community affiliated with the Australian Centre for Cyber Security (ACCS). In the case of the UNSW-NB15 dataset, the authors opt to utilize unprocessed network packets, which are generated using the highly regarded IXIA Perfectstorm program. As part of the dataset evaluation, a comprehensive range of nine attack scenarios are meticulously implemented, encompassing diverse types such as DoS, fuzzes, analysis, backdoor, generic, reconnaissance, shellcode, exploits, and worms. To provide a holistic representation of network traffic, the dataset also includes a dedicated ”Normal” class, specifically designed to capture legitimate network activity. Notably, a total of 49 network traffic features are meticulously extracted from the dataset, employing the robust Argus and Bro-IDS programs as essential analytical tools.
@@
  • WUSTL-EHMS : The WUSTL-EHMS dataset is generated using a real-time Enhanced Healthcare Monitoring System (EHMS) testbed [16]. Due to the limited availability of a dataset that integrates these biometrics, this testbed accrues both the network flow metrics and patients’ biometrics. This dataset comprises MITM attacks, spoofing and data injection. The spoofing attack merely sniffs the protocols passing through the gateway and the server, thus infringing upon the confidentiality of the patient’s data. The data injection attack is employed to dynamically modify the packets, thereby infringing upon the integrity of the data.
@@ 4.3 Data Preprocessing
During the data preprocessing phase, a systematic approach is employed to optimize the dataset for federated learning analysis. The pipeline begins with label mapping to assign binary values, enabling clear classification of instances. Following this, feature removal is performed to eliminate columns containing single values, null values, or data labeling features that could introduce bias.
The preprocessing continues with feature-type-specific transformations: Boolean features are converted to binary format, categorical features undergo either ordinal encoding or one-hot encoding based on their nature, and numerical features are standardized using StandardScaler to optimize their scale and model compatibility.
The processed data is then divided through an 80/20 train-test split. The training portion (80%) undergoes federated training data partitioning where it is distributed across multiple clients (Partition 1, Partition 2, …, Partition N) for local model training. Simultaneously, the testing portion (20%) is reserved for central server evaluation of the global model’s performance.
The flowchart depicted in Figure 5 outlines the comprehensive data preprocessing pipeline for federated learning systems, tracing the sequential transformation of raw data into partitioned datasets optimized for distributed model training and centralized evaluation.
Figure 5: Data Preprocessing and Partitioning Pipeline for Federated Learning Systems
@@ 4.4 ML Algorithm
The construction of various ANN models tailored to specific datasets involves distinctive configurations. For the UNSW-NB15 dataset, ANN are constructed with seven hidden layers, each comprising a different number of units: 150, 120, 90, 60, 30, 20, and 10 units, respectively. Conversely, the ToN-IoT dataset utilizes an ANN model with five hidden layers, incorporating 60, 40, 30, 20, and 10 units, respectively. In the case of the NSL-KDD dataset, a five-layered ANN model is established, featuring 80, 40, 30, 20, and 10 units. An ANN model is constructed for the WUSTL-EHMS dataset, comprising three hidden layers that are specified to contain 10, 20, and 40 units, respectively.
@@ 4.5 Evaluation Metrics
  • Loss : is a mathematical function used to quantify the disparity between the anticipated output of a ML model and the factual target value associated with a certain input [59]. For binary classification tasks, this difference is typically measured using the binary cross-entropy loss function, as defined in Formula (4).
@@ 4.6 Experiment Results
Figure 6: Impact of Client Pool Size on Federated Learning Convergence Using Different Datasets
@@
Figure 7: Effect of Client Participation Rate on Federated Learning Convergence Using Different Datasets
@@
Figure 8: Impact of Local Epochs on Federated Learning Convergence Using Different Datasets
@@ 4.6.1 FL Parameters Selections
To evaluate the impact of varying the number of clients on performance and communication cycles, the local epochs are fixed at 1, the fraction fit at 1, and the number of clients is adjusted to 2, 4, 8, and 12. Increasing the number of clients results in a significant reduction in communication cycles required to achieve target performance levels. This trend is evident in the Ton-IoT and NSL-KDD datasets, where the lowest number of communication cycles is achieved with 12 clients, and no significant variation is observed with 8 clients compared to other configurations, as shown in Figure 6 and Tables 3(a) and 3(c). This suggests that communication cycles stabilize with 8 or more clients. For the UNSW_NB15 dataset, the lowest number of communication cycles is achieved with 12 clients, showing a significant reduction compared to 8 clients, a trend supported by Figure 6 and Table 3(b). This aligns with the decreasing trend in communication cycles as the number of clients increases, consistent with observations from the Ton-IoT and NSL-KDD datasets. In contrast, the WUSTL-EHMS dataset exhibits an inverse trend, where increasing the number of clients leads to a higher number of communication rounds to achieve target performance, which is distinctly illustrated in Figure 6 and depicted in Table 3(d). This behavior is likely due to the limited sample size in the dataset, resulting in insufficient data distribution across clients and requiring additional communication rounds for model generalization.
@@
Increasing the fraction fit effectively reduces communication cycles while maintaining target performance levels. This trend is observed in the Ton-IoT and NSL-KDD datasets, as illustrated in Figure 7 and Tables 4(a) and 4(c). However, for the UNSW-NB15 and WUSTL-EHMS datasets, the lowest number of communication rounds is achieved with a fraction fit of 0.5, with a slight increase observed at a fraction fit of 1, a non-monotonic relationship clearly shown in Figure 7 and detailed in Tables 4(b) and 4(d). This confirms the general trend of reduced communication rounds with higher fraction fit values, though the optimal value can be dataset-dependent.
Finally, the analysis investigates the impact of the number of local epochs on performance and communication cycles. The fraction fit is set to 1, as this configuration is shown to minimize communication rounds. The number of local epochs is varied to 1, 2, 5, and 8.
A notable finding is the consistent reduction in communication cycles with an increase in local epochs while maintaining strong performance. This trend is observed across most datasets, as clearly depicted in Figure 8 and shown in Tables 5(a), 5(b), and 5(d). However, the NSL-KDD dataset achieves the optimal number of communication rounds with only 2 epochs, a deviation from the overall trend that is visually apparent in Figure 8 and illustrated in Table 5(c), though the results remain close to those obtained with 8 epochs. This further supports the overall trend of reduced communication rounds with increased local epochs, while also highlighting dataset-specific idiosyncrasies.
@@ 4.6.3 Centralized vs FL :
The results are consolidated into stacked bar plots, as depicted in Figure 13, where FL results are represented in orange, and centralized results are represented in blue. The outcomes on the NSL-KDD, TON-IOT, UNSW_NB15, and WUSTL-EHMS datasets, presented in Tables 13(a), 13(b), 13(c), and 13(d) respectively, show bars being halved, indicating comparable results between FL and centralized approaches. Consequently, training models on separate data partitions generalizes as effectively as training a global model on the entire dataset. This dual advantage of achieving comparable detection outcomes while preserving data privacy highlights the efficacy and privacy-centric nature of the FL-based IDS, positioning it as a robust and privacy-aware solution for intrusion detection.
@@ 6 DISCUSSION
The proposed solution places a paramount emphasis on safeguarding patient privacy by employing FL, which involves sharing model weights instead of actual patient data. By optimizing FL parameters, the communication rounds were significantly reduced, thereby minimizing bandwidth consumption, preventing network congestion, and ensuring the scalability of the system. The distributed nature of FL proves instrumental in positioning the IDS in close proximity to potential attack sources. This proximity enhances the system’s agility, allowing for rapid and effective detection and response to security threats. This decentralized approach strengthens the overall security posture while contributing to the rapid identification and mitigation of potential risks to patient data within the IoMT ecosystem.